# Identifying preferred routes of sharing information on social networks

### v1

@@ 3.2.2 Functional Similarity
To further investigate the presence of preference within the underlying networks of contextually similar hashtags, we focus on the behavior of individual nodes when exposed to messages of related topics. That is, we aim to determine which nodes a given node is most likely to share a received message with. A suitable metric for this purpose is the cosine similarity [newman2004detecting, newman2012communities]. Generally, the cosine similarity between two normalized vectors is given by their inner product, which quantifies the degree of alignment between the vectors in their respective vector space. When these vectors represent the sharing behavior of a specific node across two different networks, the cosine similarity reflects how similarly that node functions in those networks, hence also called the functional similarity. For the weighted, undirected network corresponding to hashtag $m_{1}$, represented by adjacency matrix $\mathbf{M_{1}}$, the retweet state vector of node $i$ is constructed as described in Eq. 9, and normalized by the norm of the space (inner product),
@@
where ${M_{1}}_{,ij}$ shows the frequency of retweets between node $i$ and node $j$ in hashtag $m_{1}$. The functional similarity of a node in the dissemination of two hashtags $m_{1}$ and $m_{2}$ with adjacency matrices $\mathbf{M_{1}}$ and $\mathbf{M_{2}}$ is then determined by
@@
Figure 4: Retweet imprints of a single node in two hashtags: In sharing two different hashtags, a node of interest (marked with gradient color) can act with partial similarity, calculated by Eq 10. The functional similarity of the node $C$ in the above example will be $S_{C}^{black,red}=\braket{p^{(black)}_{C}|p^{(red)}_{C}}\approx 0.82$. One way of measuring the overall similarity of two hashtags is to take the average of the functional similarities of all their nodes.
Similar measures based on inner product have been used for the study of retweet networks, such as favoritism of nodes [kwak2010twitter, almaas2004global]. However, favoritism is measured for a single node over one set of retweets, whereas functional similarity studies a single node over two sets of retweets (the difference might be analogized to the difference between autocorrelation and cross-correlation of series). The closer the functional similarity of a node is to unity over two hashtags, the more preferential the underlying network is.
@@ 3.2.3 Considerations
    Two hashtag datasets collected over similar time spans will rarely produce retweet network adjacency matrices of the same size; that is, the number of users involved in the sharing of each hashtag can differ significantly. However, to apply the modified weighted Jaccard index and functional similarity measures (Eqs. 8 and 10), the underlying networks must be of equal size. To assimilate the dimensions of two retweet networks, we can either expand their nodes to include the union of their users by adding isolated nodes or shrink them to sub-networks of them that include only the intersection of their users. The only difference is that the former will have additional zero terms in the summations of these measurements. The latter will eliminate the users who had only participated in the propagation of one of the two hashtags. In fact this makes our study more sensible, since our goal is to find whether a user will share similar topics from the same old route or not. So the user should participate in sharing both of them in the first place before we can even begin to study the overlap or similarity of the propagation routes.
@@ 3.3 Simulation
For the simulation, two initial underlying networks of $N=400$ were evolved under two rules: the global preference dynamic (section 2.1) and the local preference dynamic (section 2.2). For the initial network of the global preference model, $m_{0}=5$ links with weights of unity were randomly added to a weighted, undirected null network, while for the local preference model, the initial network was a fully connected network with link weights of unity. These two networks evolved independently to generate the underlying network on which we later simulate biased random walks, corresponding to the propagation of hashtags. We selected an arbitrary evolution time for each case, guided primarily by the need to obtain meaningful results in the subsequent analysis. We chose $t=300$ for the global preference model and $t=500$ for the local preference model. At each timestep, $m=3$ links were chosen to be increased by $x=10$ units. The choice of the parameter $x$ is relative to the initial weight of the links. As a null hypothesis representing the absence of sharing preferences, we also considered a third underlying network, fully connected, with homogeneous link weights. In this network, no node has any preferential priority over others.
@@
With an adiabatic assumption that the changes of the preferences of the underlying networks happens on a greater timescale than the typical time between the rise and decay of hashtags [glasgow2013hashtag], we assume the changes in individual preferences caused by the sharing of news of specific single hashtags (for example, in a time span of 7 weeks, are negligible compared to the overall time span of users’ presence, which has a median of 3-5 years [api]. With that in mind, we simulated hashtag retweets on the mentioned three networks without the retweets affecting the preferences of nodes (i.e., the weight of the links is not changed). The retweet simulation is essentially a biased random walk starting from an initial node and spreading based on each node’s preferences. In other words, the probability $p$ of the random walker jumping to a neighboring node is proportional to the weight of the links of the node it is on.
@@
Another assumption discussed in Section 2 is that hashtags with similar topics disseminate on the same underlying network. Therefore, we repeated the simulation of the biased random walks on each of the three underlying networks for a particular number of times (128 times, chosen so that 8 repetitions were accounted for the 16 corresponding hashtags). For each repetition, we added noise to the jumping probabilities from each node. We obtained the noisy jumping probabilities with a noise intensity $\eta$ from a simple mapping:
@@
This mapping is such that for $\eta=0$ the probabilities of jumping are strictly the same as the probabilities induced by the underlying network, and for $\eta=1$, the probabilities are strictly equal with no preference. Also note that the N1N-1 in the denominator accounts for a possibility of traverse to all the nodes of the network except the one the walker is standing on. For each of the simulation repetitions, $\eta$ was picked from a gaussian distribution with a mean of $0.3$ and a standard deviation of $0.15$, where the values were bounded to the interval of $[0,1]$.
To better mimic the real world, we restricted the initiating nodes of the random walks, meaning that only a certain subset of nodes started the chain of retweets by tweeting an original tweet. A quarter of the nodes were chosen as the initiating subset of the network, corresponding to the 25% of the users of X who publish 95% of the original tweets on this platform [mcclain2021behaviors]. Finally, we used a different number of total steps for the random walk of each hashtag. The number of steps for each simulation ensemble was chosen so that the distribution of the ratio of total retweets to the number of participating users of the hashtags of the real world data was satisfied.
After performing the retweet simulation as such, it is time to apply the preference existence measures mentioned in sections 3.2.1 and 3.2.2 on each pair of the hashtags both for real data (which have been halved as explained above) and their simulated counterparts separately.
@@ 4 Results and Discussion
The results of the modified weighted Jaccard index and mean functional similarity of nodes are depicted in Fig. 6. This result shows the probability distribution of each measure when applied to all the available pairs of networks in the case of real data, along with three cases of retweet simulations conducted on the three underlying networks of section 3.3. We used Sturge’s rule to determine the number of bins in the figures.
@@
(a) Probability distribution of Jaccard index of pairs of hashtags
@@
(b) Probability distribution of mean functional similarities of nodes in pairs of hashtags
@@
Figure 5: Result of measures of the existence of preference in the dissemination of tweets.
@@
The sampling errors corresponding to the three simulation cases are shown with colored areas (where the error corresponding to the case with no preference is so low that it is hardly visible).
As is seen, both the model and the retweet data span a range of values that the model with no preference cannot acquire in both the Jaccard index and mean node similarity. This means that indeed there exist sharing preferences among users in this example social network case. Moreover, the preferential models can capture the overall trend of the preference measures in the retweet data. The deficiencies in the distributions of the models compared to the data can be attributed to the models’ abstraction of the phenomena of information sharing, where it lacks to encompass the full complexity of this social-technological process, such as the presence of modular structures in the network of users [newman2006modularity], or the heterogeneity of activity rate for the sender nodes [ying2018user, vaca2014modeling].
Figure 6: Statistics of the Kolmogorov-Smirnov test for the measures of the preference of existence. The less the value of the statistics, the more the two distributions are alike.
It is worth noting that between the two measures, the average functional similarities of the models’ simulations is more comparable to the data than the results of the modified weighted Jaccard index. This is to be expected, since the latter relies on all the steps of the sharing of one hashtag, while the node similarity only studies how each individual node behaves regardless of what happens in the next steps (e.g. whether the next node will choose the more likely choice of sharing or not).
@@ 5 Conclusion
In this study, we found that the spreading of news on social networks is not random and that there are specific preferential paths for this news flow. Knowing these paths will play an important role in managing or limiting its dissemination. We introduced two mathematical models of local and global preferences to evolve networks with no preference into networks that exhibit preferential behaviors similar to those of a real-world case (e.g., retweets of hashtags on the X social platform). We measured preferential behavior using a modified weighted Jaccard index and mean node similarity. A sample of retweet data of trending hashtags confirmed the effectiveness of these two models and the preferential dissemination of news on social networks. The results of the simulated models, although differs in details from the real world data, captures the essence of the existence of preference in news dissemination seen in real world data, which significantly differs from that of a no preference model. This suggest that the mechanism of the birth of preferences for social users can follow the dynamics of the introduced models. It is important to emphasize that these preferred transmission paths are dependent on the type of news. A user’s choice of recipient may vary based on the content; for example, one contact may be preferred for political news, while another is favored for humorous content.
Future works may expand or combine global and local preferential models to better resemble a case of real world interest. The choice of the evolution time of the underlying network is another case that can be studied further and can potentially give insight into the age of the underlying network. It is also possible to study how the underlying network of two hashtags with considerably different topics differs from one another using the same measures of preference existence.
Author contributions statement
R.M. contributed to the conception of the project, conducted data analysis and simulations, and wrote the manuscript. B.A. contributed to project conception and manuscript writing. P.B. contributed to project conception and data handling. R.J. contributed to project conception, manuscript writing, and supervised the overall project.
Additional information
Competing interests
The authors declare no competing interests.
Data availability
The simulation code used in this study is available at the project’s GitHub repository. The datasets supporting this article have been uploaded on Dryad[mohammadikian2025retweet].

### v2

@@ 3.2.2 Functional Similarity
To further investigate the presence of preference within the underlying networks of contextually similar hashtags, we focus on the behavior of individual nodes when exposed to messages of related topics. That is, we aim to determine which nodes a given node is most likely to share a received message with. A suitable metric for this purpose is the cosine similarity [37, 39]. Generally, the cosine similarity between two normalized vectors is given by their inner product, which quantifies the degree of alignment between the vectors in their respective vector space. When these vectors represent the sharing behavior of a specific node across two different networks, the cosine similarity reflects how similarly that node functions in those networks, hence also called the functional similarity. For the weighted, undirected network corresponding to hashtag $m_{1}$, represented by adjacency matrix $\mathbf{M_{1}}$, the retweet state vector of node $i$ is constructed as described in Eq. 9, and normalized by the norm of the space (inner product),
@@
where ${M_{1}}_{,ij}$ shows the frequency of retweets between node $i$ and node $j$ in hashtag $m_{1}$. The functional similarity of a node in the dissemination of two hashtags $m_{1}$ and $m_{2}$ with adjacency matrices $\mathbf{M_{1}}$ and $\mathbf{M_{2}}$ is then determined by:
@@
Figure 4: Retweet imprints of a single node in two hashtags: In sharing two different hashtags, a node of interest (marked with gradient color) can act with partial similarity, calculated by Eq. 10. The functional similarity of the node $C$ in the above example will be $S_{C}^{black,red}=\braket{p^{(black)}_{C}|p^{(red)}_{C}}\approx 0.82$. One way of measuring the overall similarity of two hashtags is to take the average of the functional similarities of all their nodes.
Similar measures based on inner product have been used for the study of retweet networks, such as favoritism of nodes [28, 1]. However, favoritism is measured for a single node over one set of retweets, whereas functional similarity studies a single node over two sets of retweets (the difference might be analogized to the difference between autocorrelation and cross-correlation of series). The closer the functional similarity of a node is to unity over two hashtags, the more preferential the underlying network is.
@@ 3.2.3 Considerations
    Two hashtag datasets collected over similar time spans will rarely produce retweet network adjacency matrices of the same size; that is, the number of users involved in the sharing of each hashtag can differ significantly. However, to apply the modified weighted Jaccard index and functional similarity measures (Eqs. 8 and 10), the underlying networks must be of equal size. To assimilate the dimensions of two retweet networks, we can either expand their nodes to include the union of their users by adding isolated nodes or shrink them to sub-networks that include only the intersection of their users. The only difference is that the former will have additional zero terms in the summations of these measurements. The latter will eliminate the users who had only participated in the propagation of one of the two hashtags. In fact this makes our study more sensible, since our goal is to find whether a user will share similar topics from the same old route or not. So the user should participate in sharing both of them in the first place before we can even begin to study the overlap or similarity of the propagation routes.
@@ 3.3 Simulation
For the simulation, two initial underlying networks of $N=400$ were evolved under two rules: the global preference dynamic (Section 2.1) and the local preference dynamic (Section 2.2). For the initial network of the global preference model, $m_{0}=5$ links with weights of unity were randomly added to a weighted, undirected null network, while for the local preference model, the initial network was a fully connected network with link weights of unity. These two networks evolved independently to generate the underlying network on which we later simulate biased random walks, corresponding to the propagation of hashtags. We selected an arbitrary evolution time for each case, guided primarily by the need to obtain meaningful results in the subsequent analysis. We chose $t=300$ for the global preference model and $t=500$ for the local preference model. At each timestep, $m=3$ links were chosen to be increased by $x=10$ units. The choice of the parameter $x$ is relative to the initial weight of the links.
To benchmark the proposed global and local preference models, we implement a third network evolution model introduced by Barrat, Barthélemy, and Vespignani (BBV) [4]. In this model, new nodes attach preferentially to existing nodes according to their weighted degrees (or their strengths), which also update the existing link weights proportional to the current interaction intensity. We use the standard formulation described in [4], with simulation parameters of $N=400$, $m=3$ to match the configuration of the global and local preference simulations, and also $\delta=6$ and $w_{0}=4$. The choice of $\delta$ and $w_{0}$ was made so the receiver node’s weighted degree increases similar to the simulations of the global and local preference models (in other words, $\delta+w_{0}=x=10$, according to the BBV model. However the difference is that in the BBV model, the sender’s weighted degree would only increase $w_{0}$ units, unlike the global and local preference models, which is one of its inherent properties.
As a null hypothesis representing the absence of sharing preferences, we also considered a fourth underlying network, fully connected, with homogeneous link weights. In this network, no node has any preferential priority over others.
@@
With an adiabatic assumption that the changes in the preferences of the underlying networks happen on a greater timescale than the typical time between the rise and decay of hashtags [16], we assume the changes in individual preferences caused by the sharing of news of specific single hashtags (for example, in a time span of 7 weeks) are negligible compared to the overall time span of users’ presence, which has a median of 3-5 years [46]. With that in mind, we simulated hashtag retweets on the mentioned four networks without the retweets affecting the preferences of nodes (i.e., the weight of the links are not changed). The retweet simulation is essentially a biased random walk starting from an initial node and spreading based on each node’s preferences. In other words, the probability $p$ of the random walker jumping to a neighboring node is proportional to the weight of the links of the node it is on.
@@
Another assumption discussed in Section 2 is that hashtags with similar topics disseminate on the same underlying network. Therefore, we repeated the simulation of the biased random walks on each of the four underlying networks for a particular number of times (128 times, chosen so that 8 repetitions accounted for the 16 corresponding hashtags). For each repetition, we added noise to the jumping probabilities from each node. We obtained the noisy jumping probabilities with a noise intensity $\eta$ from a simple mapping:
@@
This mapping is such that for $\eta=0$ the probabilities of jumping are strictly the same as the probabilities induced by the underlying network, and for $\eta=1$, the probabilities are strictly equal with no preference. Also note that the N1N-1 in the denominator accounts for a possibility of traversing to all the nodes of the network except the one the walker is standing on. For each of the simulation repetitions, $\eta$ was picked from a Gaussian distribution with a mean of $0.3$ and a standard deviation of $0.15$, where the values were bounded to the interval of $[0,1]$.
To better mimic the real-world, we restricted the initiating nodes of the random walks, meaning that only a certain subset of nodes started the chain of retweets by tweeting an original tweet. A quarter of the nodes were chosen as the initiating subset of the network, corresponding to the 25% of the users of X who publish 95% of the original tweets on this platform [33]. Finally, we used a different number of total steps for the random walk of each hashtag. The number of steps for each simulation ensemble was chosen so that the distribution of the ratio of total retweets to the number of participating users of the hashtags of the real world data was satisfied.
After performing the retweet simulation as such, it is time to apply the preference existence measures mentioned in Sections 3.2.1 and 3.2.2 on each pair of the hashtags both for the empirical data and their simulated counterparts separately.
@@ 4 Results and Discussion
The results of the modified weighted Jaccard index and mean functional similarity of nodes are depicted in Fig. 5. This result shows the probability distribution of each measure when applied to all the available pairs of networks in the case of the empirical data, along with three cases of retweet simulations conducted on the four underlying networks of Section 3.3. We used Sturge’s rule to determine the number of bins in the figures.
@@
(a) Probability distribution of modified Jaccard index of pairs of hashtag retweets in data and simulations.
@@
(b) Probability distribution of mean functional similarities of nodes in pairs of hashtag retweets in data and simulations.
@@
Figure 5: Results of measures of the existence of preference in the dissemination of tweets. The No Preference and the BBV models act as the null model and benchmark model respectively.
@@
The sampling errors corresponding to the four simulation cases are shown with colored areas (where the errors corresponding to the BBV and No Preference are so low that they are hardly visible).
The deficiencies in the distributions of the proposed models against the retweet data show themselves in the p-values obtained by the Kolmogorov-Smirnov (KS) test (Table 1), calculated between the retweet data and the retweet simulation on the four models. This lack of exact fit to the data can be attributed to the two models’ abstraction of the phenomena of information sharing, where it fails to encompass the full complexity of this social-technological process, such as the presence of modular structures in the network of users [38], the heterogeneity of activity rates for the sender nodes [54, 50], or even exogenous factors affecting the retweet tendencies [2].
However, as is seen in Fig. 5, both the model and the retweet data span a range of values that the BBV and No Preference models cannot acquire in both the Jaccard index and functional similarity. This shows that, firstly, indeed there exist sharing preferences among users in this example social network case which the No Preference model cannot capture for obvious reasons, and secondly, the mechanism that generates the sharing preference in our proposed models works better at containing the spectrum of the results in the empirical data than the benchmark model, the BBV. Comparison of distributions in Fig. 5(a) and 5(b) is quantified in Fig. 6, where the Jensen-Shannon Divergence (JSD) [32] was calculated between the empirical retweet data and the four models’ retweet simulations, which measures how distinguishable two distributions are by averaging how far each distribution is from their shared midpoint distribution. This measure is bounded between zero (distributions are exactly the same) and $ln2\approx 0.693$ (distributions are completely different). The JSD reports an overall information difference between two distributions, whereas the KS statistics indicates the largest local discrepancy between them. Thus for a quantitative comparison of two distributions which were not expected to fit exactly as discussed before, the JSD is a more descriptive measure.
It is worth noting that while the modified Jaccard index relies on all steps of the sharing of one hashtag and how all the nodes involved have behaved, the functional similarity is affected only by each individual node’s preference, regardless of how the introduced stochasticity in the retweet simulation along with the topology of the underlying network might affect the whole path, thus rendering a better measure of the existence of individual preferences. In Fig. 6, it is seen that the modified Jaccard index is closer to the real-life data than its mean functional similarity measure for the BBV model, considering the errors. In contrast the global and local preference models yield reverse results, showing that their mechanisms have been more successful in producing individual sharing preferences such as the real life social networks.
@@ 5 Conclusion
In this study, we found that the spreading of news on social networks is not random and that there are specific preferential paths for this news flow. Knowing these paths will play an important role in managing or limiting its dissemination. We introduced two mathematical models of global and local preferences to evolve networks with no preference into networks that exhibit preferential behaviors similar to those of a real-world case (e.g., retweets of hashtags on the X social platform). We measured preferential behavior using a modified weighted Jaccard index and mean functional similarity. A sample of retweet data of trending hashtags confirmed the effectiveness of these two models and the preferential dissemination of news on social networks, compared to a benchmark and a null model. The results of the simulated models, although they differ in details from the real-world data, capture the essence of the existence of preference in news dissemination seen in real-world data. This suggests that the mechanism of the birth of preferences for social users can follow the dynamics of the introduced models. It is important to emphasize that these preferred transmission paths are dependent on the type of news. A user’s choice of recipient may vary based on the content; for example, one contact may be preferred for political news, while another is favored for humorous content.
Future works may expand or combine global and local preferential models to better resemble a case of real-world interest. The choice of the evolution time of the underlying network is another case that can be studied further and can potentially give insight into the age of the underlying network, which directly connects to temporal analysis over larger time frames. It is also possible to study how the underlying network of two hashtags with considerably different topics differs from one another using the same measures of preference existence.
Modified Jaccard Index
Simulated Model Statistic p-value
Global 0.469 $10^{-15}$
Local 0.156 $5.547\times 10^{-5}$
BBV 0.142 $7.086\times 10^{5}$
No Preference 0.913 $10^{-15}$
Mean Functional Similarity
Simulated Model Statistic p-value
Global 0.179 $1.96\times 10^{-6}$
Local 0.298 $10^{-15}$
BBV 0.350 $10^{-15}$
No Preference 0.922 $10^{-15}$
Table 1: KS statistics and corresponding p-values of the comparison of the empirical data from X vs. the retweet simulation based on the two proposed models (global and local), the benchmark model (BBV) and the null model (No Preference), calculated for the two preference measures. P-values less than the machine’s precision are presented by $10^{-15}$.
Figure 6: JSD for the results of the preference measures between the empirical data from X and the four simulated models. The JSD measure is bounded in the $0,\ln 2]$ interval, with zero indicating perfect fit between two distributions and $\ln 2$ indicating maximal difference.
Acknowledgment
We would like to thank Parham Moradi for his valuable assistance with data collection.